Identify whether each of the followings is permutation or combination and find the

answers.

(i) lsha wants to invest 12 million in three projects. She decided to invest 4 million in

each project and if seven shortlisted projects are available, in how many possible

ways she can invest her money?

(ii) Instead of equal allocation, she decided to invest 3million, 4 million and 5 million for

each project. How many possible arrangements are available for her investment

decision?



1
Expert's answer
2021-12-01T13:48:10-0500

(i) The order does not matter. We use combination



"C(7,3)=\dfrac{7!}{3!(7-3)!}=35"

(ii) The order matters. We use permutation



"P(7,3)=\dfrac{7!}{(7-3)!}=210"